OpenStudy (anonymous):

If you invest $25,000 in an account that gets 12% annual interest compounded quarterly, how much would you have in 10 years.

OpenStudy (cwrw238):

start again!: after 10 years amount = 25000(1 + (0.12)/4)^(10*4)

OpenStudy (anonymous):

okay.. so what would that be ?

OpenStudy (cwrw238):

this simplifies to 25000(1.03)^40 - you can compute that on your calculator

OpenStudy (cwrw238):

you should get 81,550.94 dollars

OpenStudy (cwrw238):

the general formular is A = P(1 + N/r)^(Nn) where P = principal (starting amount), r = annual rate, N = number of times per year the interest is compounded, and n = number of years

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

okay so whats does it look like all finished?>

OpenStudy (cwrw238):

well thats it the amount after 10 years is 81,550.94 - pretty good rate of interest